# GT2-GS: Geometry-aware Texture Transfer for Gaussian Splatting

### v1

@@ Abstract.
Transferring 2D textures to 3D modalities is of great significance for improving the efficiency of multimedia content creation. Existing approaches have rarely focused on transferring image textures onto 3D representations. 3D style transfer methods are capable of transferring abstract artistic styles to 3D scenes. However, these methods often overlook the geometric information of the scene, which makes it challenging to achieve high-quality 3D texture transfer results. In this paper, we present GT2-GS, a geometry-aware texture transfer framework for gaussian splitting. From the perspective of matching texture features with geometric information in rendered views, we identify the issue of insufficient texture features and propose a geometry-aware texture augmentation module to expand the texture feature set. Moreover, a geometry-consistent texture loss is proposed to optimize texture features into the scene representation. This loss function incorporates both camera pose and 3D geometric information of the scene, enabling controllable texture-oriented appearance editing. Finally, a geometry preservation strategy is introduced. By alternating between the texture transfer and geometry correction stages over multiple iterations, this strategy achieves a balance between learning texture features and preserving geometric integrity. Extensive experiments demonstrate the effectiveness and controllability of our method. Through geometric awareness, our approach achieves texture transfer results that better align with human visual perception. Our homepage is available at this https url.
Texture Transfer, 3D Gaussian Splatting, Geometry-aware
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Rendering
@@
Figure 1. Previous methods struggle to accurately transfer complex texture appearances to the scene. In contrast, our GT2-GS framework injects geometric information into the scene during texture transfer, enabling geometry-aware texture transfer. Our method enables continuous and controllable texture transfer, and is capable of producing perspective effects as illustrated by the blue boxes.
1. Introduction
Texture transfer (Ashikhmin, 2003; Efros and Freeman, 2023; Wang et al., 2022) aims to transfer the texture features of a reference image to a target image while preserving the structural and semantic content. Early texture transfer algorithms (Ashikhmin, 2003; Lee et al., 2010; Hertzmann et al., 2023; Efros and Freeman, 2023) were based on non-parametric texture synthesis and were mainly applied to tasks such as image inpainting and data augmentation. With the advancement of deep learning, a variety of network-based texture transfer methods (Wang et al., 2022; Chen et al., 2022; Cao et al., 2023; Pu et al., 2024) have emerged. These methods enable texture transfer by incorporating semantic information, allowing for artistic content creation.
With the rapid development of fields such as virtual reality, robotics, film, and gaming, the demand for high-quality 3D content has grown significantly. Texture is an essential component of various graphics pipelines and plays a crucial role in producing high-quality 3D assets. Traditionally, texture creation requires skilled artists to design them manually, which is both costly and inefficient. Recently, the emergence of neural radiance fields (NeRF) (Mildenhall et al., 2020) and 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) has greatly improved the efficiency of 3D content creation. Building upon such 3D representations, it is promising to develop 3D texture transfer method that directly transfers texture features from reference images to 3D representations. This approach bypasses the labor-intensive UV mapping process, significantly reducing both the cost and technical barriers of 3D content creation. Due to the challenge of multi-view consistency, 2D texture transfer methods (Ashikhmin, 2003; Efros and Freeman, 2023; Wang et al., 2022) cannot be directly applied to 3D scenes. Some diffusion-based approaches (Cao et al., 2023; Mitchel et al., 2024; Yeh et al., 2024; Xie et al., 2024) generate high-quality texture maps for 3D meshes. However, these methods modify the appearance of 3D objects via texture maps rather than directly transferring texture features into 3D representations. Moreover, they typically operate at the object level, making it difficult to generate textures for entire scenes. Therefore, how to achieve efficient 3D texture transfer remains an open and worthwhile research question.
Existing 3D scene stylization methods (Zhang et al., 2022; Liu et al., 2023, 2024; Zhang et al., 2024a; Liu et al., 2025) focus on transferring the artistic styles of abstract reference images to 3D scenes. We attempt to use texture images as references and apply state-of-the-art optimization-based stylization methods to transfer their texture features to 3D scenes. However, as shown in the first row on the left side of Fig. 1, the texture transfer results of this approach are suboptimal. Through analysis, we identify the core reason why existing 3D stylization methods cannot be directly applied to texture transfer. Texture is closely tied to geometry. 3D texture transfer needs to be harmonious with both the 3D geometry and the texture features of the reference image, requiring consistency between texture and geometry to ensure strict multi-view consistency. However, existing 3D stylization methods overlook the geometry of the scene. Firstly, 3D scenes are optimized via gradient descent using losses computed between rendered feature maps and texture features. Since texture images typically contain only a single type of texture, applying highly repetitive texture features across diverse rendered views is suboptimal. This fails to align with the rich geometric information in the rendered images, such as perspective geometry, edges, contours, and shapes. Moreover, current style transfer loss functions are computed independently across views, without considering the 3D geometric relationships between different viewpoints. This limits the ability to achieve spatially consistent and controllable texture transfer.
In this paper, we propose a novel framework GT2-GS for achieving geometry-aware texture transfer. Our insight is to integrate comprehensive geometric information into the texture transfer process to achieve controllable results that align with human visual perception. To enable geometry-aware texture transfer, we first propose a geometry-aware texture augmentation module. This module enhances texture features through scaling and rotation operations based on geometric information from the rendered image, such as perspective, edges, and contours. In addtion, a geometry-consistent texture Loss is proposed for 3D texture transfer. This loss uses the camera poses and depth geometry between a prior view and the current optimization view, achieving scene appearance editing with controllable texture orientation. To balance texture learning and geometry preservation, we further introduce a Geometry Preservation Strategy. This strategy alternates the optimization of Gaussian parameters between a texture transfer stage and a geometry correction stage, with multiple iterations performed. In this process, the geometry correction stage utilizes content images with ground truth for optimization, ensuring the accurate refinement of scene geometry. Extensive experiments demonstrate that our proposed framework achieves high-quality texture transfer results.
@@ 1. Introduction
  • To the best of our knowledge, we are the first to incorporate geometric information into 3D texture transfer. The proposed framework enables direction-controllable and high-quality texture transfer results.
  • Geometry-aware texture Augmentation module and geometry-consistent texture loss are proposed to incorporate the geometric information of the scene during the 3d scene texture transfer. Moreover, we introduced a geometric preservation strategy to maintain the correct geometric information of the scene.
@@
2. Related Work
2.1. Texture Transfer
@@
Early texture transfer algorithms (Ashikhmin, 2003; Lee et al., 2010; Hertzmann et al., 2023; Efros and Freeman, 2023) were based on non-parametric texture synthesis. While preserving the structural information of the target image to achieve texture transfer effects. Gatys et al. (Gatys et al., 2016) pioneered the use of convolutional neural networks (CNNs) (Simonyan and Zisserman, 2014) for image style transfer, achieving impressive results. Subsequently, a variety of deep learning-based style transfer methods (Chen and Schmidt, 2016; Li and Wand, 2016; Huang and Belongie, 2017; Huo et al., 2021) have been proposed. These methods are capable of transferring abstract artistic elements such as color distribution, brushstroke characteristics, and overall composition, rather than focusing solely on texture transfer. Leveraging the powerful feature extraction capabilities of networks such as CNNs and Transformers (Vaswani et al., 2017), some deep learning-based texture transfer methods (Wang et al., 2022; Chen et al., 2022; Pu et al., 2024) can transfer local textures from the reference image to the target image based on high-level semantic correspondence. With the development of generative models (Goodfellow et al., 2014; Ho et al., 2020), some methods (Lyu et al., 2021; Cao et al., 2023; Mitchel et al., 2024; Yeh et al., 2024; Xie et al., 2024) have adopted texture images as guidance conditions for texture transfer. Notably, high-quality 3D texture maps can be directly generated using diffusion-based methods (Cao et al., 2023; Mitchel et al., 2024; Yeh et al., 2024; Xie et al., 2024). However, these methods require precise computation on existing meshes and are difficult to extend to complex and large-scale scenes. In this work, we explore how to directly transfer texture features into 3D representations.
@@
Figure 2. Overview of GT2-GS. The overall pipeline is illustrated in (a). The input to our proposed framework includes the scene Gaussians, a content image, and a texture image. The texture image is transformed into a texture feature set through the GTA module, as shown in (b). This feature set is then used to construct the target feature map in the GT Loss, illustrated in (c). After embedding the parameter $c^{g}$, the Gaussians are optimized using the two-stage iterative strategy GPS, as shown in (d). As a result, the texture features are seamlessly integrated into the Gaussian representation.
2.2. 3D Style Transfer
@@
3D style transfer aims to transfer the style from 2D images to the appearance of 3D scenes, while maintaining the scene’s content and multi-view consistency. Early research in 3D scene style transfer focused on explicit representations such as point clouds (Cao et al., 2020; Huang et al., 2021) ,meshes (Höllein et al., 2022; Yin et al., 2021), and voxels (Guo et al., 2021; Klehm et al., 2014). Zhang et al. (Zhang et al., 2022) explored the potential of NeRF in 3D stylization tasks. By optimizing the scene representation with the proposed 3d style transfer loss, it achieved pleasing visual effects. Subsequently, a variety of NeRF-based 3D style transfer methods (Huang et al., 2022; Pang et al., 2023; Liu et al., 2023; Zhang et al., 2024b; Jung et al., 2024) have emerged. Recently, the emergence of 3DGS (Kerbl et al., 2023) has brought new possibilities for scene stylization. It features fast training speed, high rendering quality, and efficient performance. Some works (Liu et al., 2024; Zhang et al., 2024a; Mei et al., 2024; Liu et al., 2025) have already explored the potential of stylization under this representation. Currently, 3DGS-based scene stylization methods can be categorized into forward-feed approaches (Liu et al., 2024) and optimization-based approaches (Zhang et al., 2024a; Mei et al., 2024; Liu et al., 2025). Forward-feed methods enable zero-shot style transfer. For instance, StyleGaussian (Liu et al., 2024) embeds VGG features into Gaussians and leverages pre-trained scene representations combined with AdaIN (Huang and Belongie, 2017) to achieve real-time stylization. On the other hand, optimization-based methods attain higher-quality style transfer results. Some works (Zhang et al., 2024a; Liu et al., 2025) further exploit the explicit characteristic of Gaussian representations to enable region-controllable scene stylization.
However, existing 3D stylization methods do not incorporate geometric information during the stylization process, making them unsuitable for texture images. Our work focuses on texture transfer. By introducing geometric information, we achieve high-quality results that align with human visual perception.
3. Preliminaries
3.1. 3D Gaussian Splatting
@@ 3. Preliminaries
3D Gaussian Splatting (Kerbl et al., 2023) is a novel explicit 3D representation method. It uses a set of parameterized Gaussians $G=\{g_{i}\}$ to represent 3D scenes. The paramaters of each Gaussian $g_{i}$ include a mean vector $\mu_{i}$ representing its center position, a covariance matrix $\Sigma_{i}$ describing its shape, a opacity parameter $\sigma_{i}$ and a color parameter $c_{i}$ which is represented as spherical harmonic coefficients. Among these, the covariance matrix $\Sigma_{i}$ is decomposed into rotation parameters $r_{i}$ and scaling parameters $s_{i}$ to effectively maintain its positive semi-definite property during the optimization process. The optimization parameters of Gaussian $g_{i}$ are actually represented as $g_{i}=\{\mu_{i},r_{i},s_{i},\sigma_{i},c_{i}\}$. The 3D Gaussians are efficiently rendered through a fast differentiable rasterization (Lassner and Zollhofer, 2021). Specifically, the Gaussians are grouped into tiles and sorted by depth. The color $C$ of a pixel is computed through $\alpha$-blending:
(1) $C=\sum_{i\in{N}}T_{i}\alpha_{i}c_{i},T_{i}=\prod_{j=1}^{i-1}(1-\alpha_{j}),$
@@
where $T_{i}$ is the transmittance and $\alpha_{i}$ is the alpha-compositing weight for the $i$-th Gaussian. The initial Gaussians are initialized from a sparse point cloud obtained via Structure-from-Motion (SfM) (Schonberger and Frahm, 2016) using the input multi-view images. During optimization, 3DGS adopts an adaptive densification strategy to control the distribution of Gaussians. Specifically, this strategy performs cloning or splitting of Gaussians based on their positional gradients and sizes. Moreover, it periodically removes Gaussians with low opacity or large geometric extent.
3.2. Optimization for 3D stylization
@@
Optimization for 3D stylization requires balancing both style transfer effects and the preservation of original content information. By computing the loss on the feature maps of rendered views, the learnable parameters in the scene representation can be optimized. Existing optimization-based 3D style transfer methods (Zhang et al., 2022; Pang et al., 2023; Zhang et al., 2024b, a) typically adopt the nearest neighbor feature matching (NNFM) loss (Zhang et al., 2022) as their style loss function. It matches the rendered features with the nearest neighbor features in the style feature set and minimizes the cosine distance between them. Specifically, a random viewpoint is selected to obtain the rendered image $I_{r}$. The same feature extractor (e.g., VGG (Simonyan and Zisserman, 2014)) is used to extract the feature maps $F_{r}$ and $F_{s}$ from $I_{r}$ and the style image $I_{s}$, respectively. Let $F_{r}(i,j)$ denote the feature vector at the pixel location $(i,j)$ of the rendered feature map $F_{r}$, the NNFM loss can be expressed as
(2) $L_{style}=\frac{1}{N}\sum_{i,j}\min_{i^{\prime},j^{\prime}}dist(F_{r}(i,j),F_{s}(i^{\prime},j^{\prime})),$
@@
where $N$ is the number of pixels in $F_{r}$, $dist(a,b)$ is the cosine distance between two feature vectors $a$ and $b$.
The content loss $L_{content}$ is the perceptual loss, which is computed using the $L_{2}$ loss between the rendered feature map $F_{r}$ and the content feature map $F_{c}$. The total stylization loss function consists of the style loss and the content loss, expressed as
(3) $L=L_{style}+\lambda L_{content},$
where $\lambda$ is the scaling hyperparameter. By adjusting the scaling hyperparameter $\lambda$, the degree of stylization can be controlled.
4. Method
4.1. GT2-GS FrameWork Overview
In this section, we provide a detailed overview of the proposed GT2-GS framework. During the preprocessing stage, we extract a set of texture features from the input texture image using the proposed Geometry-aware Texture Augmentation (GTA) module. Moreover, Geometry-consistent Texture Loss (GT Loss) is proposed to enable controllable texture transfer. It is computed using the texture features extracted by the GTA module and the prior information from the previous viewpoints. For the input Gaussian, an additional color parameter $c^{g}$ is embedded and initialized using the current color parameter $c$. During 3D texture transfer, we proposed Geometry Preservation Strategy (GPS), which updates Gaussian parameters through multiple iterative rounds. In each round, the color parameter is first switched to $c$, and a texture transfer stage is performed. After Gaussian densification, a geometry correction stage is conducted with the color parameter switched to $c^{g}$. After several rounds of iteration, the texture features are effectively transferred to the scene representation. The pipeline is shown in Fig.  2.
4.2. Geometry-aware Texture Augmentation
The scene learns texture features through rendered images, and the scale of the texture on the object’s surface in the rendered view is related to the distance between the object and the camera. This relationship adheres to the perspective geometry principle. Simultaneously, multi-view rendered images contain rich 2D geometric information such as edges, contours, and shapes. However, texture images exhibit a high degree of repetitive patterns. The texture features extracted using the VGG feature extractor exhibit a homogenization phenomenon. The rendered feature maps fail to align with the desired texture features, thereby impairing the scene’s ability to learn the target texture appearance. To address these issues, we augment the texture features by incorporating geometric information.
To obtain multi-scale texture features that align with the perspective geometric effect, we perform texture mapping on the initial texture image by incorporating the depth map rendered from the original scene. The depth map can be obtained using a method similar to rendering color. Specifically, the depth value $D$ of a pixel is computed by blending $N_{p}$ ordered Gaussians that overlap the pixel as
(4) $D=\sum_{i\in{N_{p}}}T_{i}\alpha_{i}d_{i},$
where $d_{i}$ is the depth for $i$-th Gaussian. Subsequently, the texture image can be scaled according to the depth values. Given the excessive number of distinct depth values for all pixels in the depth map, we sort the values and discretize them into $K$ groups based on the depth intervals. Using the lowest group depth value $Z_{1}$ as the reference, the corresponding scaling factor $s$ for each group is obtained through $s_{k}=Z_{1}/Z_{k}$. The depth value $Z_{1}$ of the lowest group is determined by the upper bound of the depth interval to avoid division-by-zero issues. Meanwhile, the depth values $Z_{k}$ of the remaining groups are determined by the lower bound of their respective intervals. These scaling factors are then converted into scaling matrices $M_{s}$, which are utilized for multi-scale texture mapping by equation as $I_{scale}=M_{s}I_{s}$.
Edges, contours, and shapes, as 2D geometric features, inherently possess directional attributes. The directional information of the entire rendered image can encompass a wide range of angles. However, homogeneous texture images are insufficient to represent geometric information across all directions, particularly for textures such as lines and arrows. To address this, we apply the rotation matrix $M_{r}$ to extend the multi-scale texture in various directions $\Phi$. The complete geometric-aware texture augmentation process can be formulated as
(5) $I_{aug}^{k,\theta}={M_{r}^{\theta}}{M_{s}^{k}}I_{s},$
where $I_{aug}$ is the image obtained through augmentation, $k\in K$ is the index of depth group, and $\theta\in\Phi$ represents the rotation angle applied to the texture image.
Through the aforementioned steps, we have obtained a series of texture images based on geometric information. Although the texture feature set is extensive for scene texture transfer, the excessive number of features within the set adversely impacts the computational efficiency of the texture transfer process. Given the high homogeneity of texture features, we apply center cropping to the extracted feature maps to reduce redundant texture features. Each feature in the texture feature set is defined as $\{f_{k,\theta},s_{k},\theta\}$. The scaling and rotation parameters are retained for subsequent computations.
@@ 4.2. Geometry-aware Texture Augmentation
Figure 3. Differences in perspective. The same textured region (highlighted in red) on the 3D object exhibits different texture orientations under varying viewpoints.
4.3. Geometry-consistent Texture Loss
With the GTA module, we have already obtained sufficient texture features. However, for initial texture images with the same texture patterns but differing directions, the texture feature sets derived from geometry-aware texture enhancement are approximately identical. If the previous NNFM loss is directly used for texture transfer, it becomes impossible to control the texture orientation learned by the scene. To address the issue of controllability, we incorporate the 3D geometric relationships between different camera viewpoints into the calculation of the loss function.
GT Loss is proposed to enable geometry-consistent texture transfer. Specifically, we construct a per-pixel matched target feature map $F_{t}$ for the rendered feature map $F_{r}$. Texture transfer is achieved by minimizing the cosine similarity between the corresponding feature vectors of $F_{r}$ and $F_{t}$. To specify the texture orientation for a given viewpoint, we enforce geometric consistency through inter-view geometric relationships. Specifically, except for the viewpoint with the designated texture orientation, all other viewpoints consider the results from the previous viewpoint when constructing the target feature map. We achieve prior information transmission by using geometric information from the depth map $I_{d}^{v}$ of the current view and intrinsic and extrinsic camera parameters. Through this information, we derive a prior information transmission matrix $M_{p}^{v-1,v}$ via the projection process from the camera pose of current view $\varphi_{v}$ to 3d space to camera pose $\varphi_{v-1}$ in prior view. The computation process of this matrix is formulated as
(6) $M_{p}^{v,v-1}=K_{v-1}[R_{v-1}|T_{v-1}]{[R_{v}|T_{v}]^{-1}}{K_{v}^{-1}},$
@@ 4.3. Geometry-consistent Texture Loss
where $K_{v}$ and $[R_{v}|T_{v}]$ represent the intrinsic matrix and world-to-camera extrinsic parameters of the $v$-th viewpoint, respectively. Through the information transfer matrix, the current viewpoint can sample the prior feature $f_{k^{\prime},\theta^{\prime}}$ in the screen coordinate system of the prior viewpoint. Considering occlusion relationships, we filter the projected points using the depth map $I_{d}^{v-1}$ of the prior viewpoint. We denote $X$ as a binary mask indicating the positions in the current view’s feature map where prior information can be obtained. Specifically, $X(i,j)=1$ indicates that prior information is available at position $(i,j)$, while $X(i,j)=0$ means it is not.

### v3

@@ Abstract
Transferring 2D textures onto complex 3D scenes plays a vital role in enhancing the efficiency and controllability of 3D multimedia content creation. However, existing 3D style transfer methods primarily focus on transferring abstract artistic styles to 3D scenes. These methods often overlook the geometric information of the scene, which makes it challenging to achieve high-quality 3D texture transfer results. In this paper, we present GT2-GS, a geometry-aware texture transfer framework for gaussian splatting. First, we propose a geometry-aware texture transfer loss that enables view-consistent texture transfer by leveraging prior view-dependent feature information and texture features augmented with additional geometric parameters. Moreover, an adaptive fine-grained control module is proposed to address the degradation of scene information caused by low-granularity texture features. Finally, a geometry preservation branch is introduced. This branch refines the geometric parameters using additionally bound Gaussian color priors, thereby decoupling the optimization objectives of appearance and geometry. Extensive experiments demonstrate the effectiveness and controllability of our method. Through geometric awareness, our approach achieves texture transfer results that better align with human visual perception. Our homepage is available at https://vpx-ecnu.github.io/GT2-GS-website.
Introduction
3D style transfer (zhang2022arf; zhang2024stylizedgs) aims to transfer the stylistic elements of a reference image onto a 3D scene while preserving the scene’s original structural and semantic information. With the rapid development of fields such as virtual reality, robotics, film, and gaming, the demand for high-quality 3D content has increased significantly. 3D stylization techniques offer a promising solution by accelerating the creation of 3D content, particularly in complex 3D scene environments.
Recently, the emergence of Neural Radiance Fields (NeRF) (mildenhall2020nerf) and 3D Gaussian Splatting (3DGS) (kerbl3Dgaussians) has significantly advanced the field of 3D stylization. NeRF-based stylization methods (zhang2022arf; zhang2023ref; nguyen2022snerf), leveraging the advantages of implicit representations, allow for the decoupled optimization of appearance and geometry. In contrast, 3DGS-based approaches (galerne2025sgsst; liu2025abc; liu2024stylegaussian) offer benefits such as explicit editability and real-time rendering. 3D stylization methods can be broadly categorized into feed-forward and optimization-based approaches, depending on whether they support zero-shot style transfer. While feed-forward methods (liu2023stylerf; huang2022stylizednerf; liu2024stylegaussian) enable zero-shot stylization, they often exhibit lower rendering quality. In contrast, optimization-based methods (zhang2022arf; liu2025abc; zhang2024stylizedgs; galerne2025sgsst) typically produce higher-fidelity stylization results. Despite these differences, both types of methods commonly define optimization objectives within the VGG feature space during training.
@@ Introduction
Figure 1: Previous state-of-the-art methods struggle to accurately transfer complex textures to 3D scene. In contrast, our GT2-GS framework incorporates geometric information into the optimization process, enabling geometry-aware texture transfer.
Texture is a crucial component of various graphics pipelines and plays a vital role in producing high-quality 3D assets. Compared to abstract artistic styles, texture elements are more controllable for users. However, as illustrated in Fig. 1, existing methods struggle to accurately transfer the texture of the reference image onto the 3D scene. We analyze the limitations of current approaches from the perspective of the scene optimization process. First, during scene optimization, multi-view objectives should preserve correct geometric relationships. For example, using content images captured from the same scene as ground truth. However, existing style transfer methods define their objectives independently across views, without considering the rich geometric structure within the scene or the geometric consistency across viewpoints. Moreover, we observe a fine-grained mismatch between the VGG feature space and the pixel space. Commonly used style transfer losses fail to account for this mismatch, which leads to regions with high pixel-level information density being easily disrupted by coarse-grained feature representations. Addressing these issues is crucial for achieving high-quality texture transfer.
In this paper, we propose a novel framework GT2-GS for achieving geometry-aware texture transfer. Our method explicitly accounts for geometric information and fine-grained discrepancies, enabling texture transfer results that better align with human visual perception. We first propose a geometry-aware texture transfer loss, which is built upon texture features augmented with additional geometric parameters and incorporates cross-view geometric information to ensure consistent and controllable texture transfer. In addition, an adaptive fine-grained control module is proposed. It adaptively adjusts the strength of texture learning based on the information density of different pixel regions, thereby mitigating the fine-grained discrepancy between texture features and pixel-level representations. To address the coupling between Gaussian geometry and color parameters, we further introduce a geometry preservation branch. We bind additional color parameters to the Gaussians and optimize them using the content image as ground truth, in order to accurately refine the scene geometry. Extensive experiments demonstrate that our proposed framework achieves high-quality texture transfer results.
@@
  • We propose a geometry-aware texture transfer framework for general 3DGS scenes. The proposed framework enables controllable and high-quality texture transfer results.
  • Our Geometry-aware Texture Transfer Loss and Adaptive Fine-Grained Control Module respectively account for geometric information and fine-grained discrepancies between texture features and pixels, enabling high-quality texture transfer. Moreover, the Geometry Preservation Branch provides a novel approach to preserving geometry during appearance editing.
@@
Related Work
Texture Transfer
@@
Early texture transfer algorithms (ashikhmin2003fast; lee2010directional; hertzmann2023image; efros2023image) were based on non-parametric texture synthesis. These methods preserve the structural information of the target image to achieve texture transfer effects. Gatys et al. (gatys2016image) pioneered the use of convolutional neural networks (CNNs) (simonyan2014very) for image style transfer, achieving impressive results. Subsequently, a variety of deep learning-based style transfer methods (chen2016fast; li2016combining; huang2017arbitrary; huo2021manifold) have been proposed. These methods are capable of transferring abstract artistic elements such as color distribution, brushstroke characteristics, and overall composition, rather than focusing solely on texture transfer. Leveraging the powerful feature extraction capabilities of networks such as CNNs and Transformers (vaswani2017attention), some deep learning-based texture transfer methods (wang2022texture; chen2022auv; pu2024dynamic) can transfer local textures from the reference image to the target image based on high-level semantic correspondence. In this work, we explore how to directly transfer texture features into 3D representations.
@@
Figure 2: Overview of GT2-GS. The overall pipeline is illustrated in (a). The input to our proposed framework includes the scene Gaussians, content images, and a texture image. The texture image is transformed into a texture feature set, which is used to construct the target feature map in the GT2 Loss, as shown in (b). We extract pixel-wise information from the input and transform it into an adaptive weight matrix to control texture learning, as illustrated in (c). After embedding the parameter $c^{g}$, the Gaussians are optimized using two branches. The additional geometry preservation branch is shown in (d). As a result, the texture features are integrated into the Gaussian representation.
3D Style Transfer
@@
3D style transfer aims to transfer the style from 2D images to the appearance of 3D scenes, while maintaining the scene’s content and multi-view consistency. Early research in 3D scene style transfer focused on explicit representations such as point clouds (cao2020psnet; huang2021learning), meshes (hollein2022stylemesh; yin20213dstylenet), and voxels (guo2021volumetric; klehm2014property). Zhang et al. (zhang2022arf) explored the potential of NeRF in 3D stylization tasks. By optimizing the scene representation with the proposed 3D style transfer loss, it achieved pleasing visual effects. Subsequently, a variety of NeRF-based 3D style transfer methods (huang2022stylizednerf; pang2023locally; liu2023stylerf; zhang2024coarf; jung2024geometry) have emerged. Recently, the emergence of 3DGS (kerbl3Dgaussians) has brought new possibilities for scene stylization. It features fast training speed, high rendering quality, and efficient performance. Some works (liu2024stylegaussian; zhang2024stylizedgs; mei2024regs; liu2025abc) have already explored the potential of stylization under this representation. Currently, 3DGS-based scene stylization methods can be categorized into feed-forward approaches (liu2024stylegaussian) and optimization-based approaches (zhang2024stylizedgs; mei2024regs; liu2025abc). Feed-forward methods enable zero-shot style transfer. For instance, StyleGaussian (liu2024stylegaussian) embeds VGG features into Gaussians and leverages pre-trained scene representations combined with AdaIN (huang2017arbitrary) to achieve real-time stylization. On the other hand, optimization-based methods attain higher-quality style transfer results. Some works (zhang2024stylizedgs; liu2025abc) further exploit the explicit characteristic of Gaussian representations to enable region-controllable scene stylization. However, existing methods do not incorporate geometric information during the stylization process, making them unsuitable for texture images. Meanwhile, these methods overlook the fine-grained discrepancies between features and pixels. By incorporating geometric information and adaptive fine-grained control, our method is able to achieve high-quality texture transfer results.
Preliminaries
3D Gaussian Splatting
@@
3D Gaussian Splatting (kerbl3Dgaussians) is a novel explicit 3D representation method. It uses a set of parameterized Gaussians $G=\{g_{i}\}$ to represent 3D scenes. The parameters of each Gaussian $g_{i}$ include a mean vector $\mu_{i}$ representing its center position, a covariance matrix $\Sigma_{i}$ describing its shape, an opacity parameter $\sigma_{i}$ and a color parameter $c_{i}$, which is represented as spherical harmonic coefficients. Among these, the covariance matrix $\Sigma_{i}$ is decomposed into rotation parameters $r_{i}$ and scaling parameters $s_{i}$ to effectively maintain its positive semi-definite property during the optimization process. The optimization parameters of Gaussian $g_{i}$ are actually represented as $g_{i}=\{\mu_{i},r_{i},s_{i},\sigma_{i},c_{i}\}$. The 3D Gaussians are efficiently rendered through a fast differentiable rasterization (lassner2021pulsar). Specifically, the Gaussians are grouped into tiles and sorted by depth. The color $C$ of a pixel is computed through $\alpha$-blending:
$C=\sum_{i\in N}T_{i}\alpha_{i}c_{i},T_{i}=\prod_{j=1}^{i-1}(1-\alpha_{j}),$ (1)
where $T_{i}$ is the transmittance and $\alpha_{i}$ is the alpha-compositing weight for the $i$-th Gaussian. The depth value $D$ can be obtained using a method similar to rendering color,
$D=\sum_{i\in N}T_{i}\alpha_{i}d_{i},$ (2)
@@
where $d_{i}$ is the depth for $i$-th Gaussian. During optimization, 3DGS adopts an adaptive densification strategy to control the distribution of Gaussians. Specifically, this strategy performs cloning or splitting of Gaussians based on their positional gradients and sizes.
Style Transfer Loss
@@
By computing the loss on the feature maps of rendered views, the learnable parameters in the scene representation can be optimized. Existing optimization-based 3D style transfer methods (zhang2022arf; pang2023locally; zhang2024coarf; zhang2024stylizedgs) typically adopt the nearest neighbor feature matching (NNFM) loss (zhang2022arf) as their style loss function. It matches the rendered features with the nearest neighbor features in the style feature set and minimizes the cosine distance between them. Specifically, a random viewpoint is selected to obtain the rendered image $I_{r}$. The same feature extractor (e.g., VGG (simonyan2014very)) is used to extract the feature maps $F_{r}$ and $F_{s}$ from $I_{r}$ and the style image $I_{s}$, respectively. Let $F_{r}(i,j)$ denote the feature vector at the pixel location $(i,j)$ of the rendered feature map $F_{r}$, the NNFM loss can be expressed as
$L_{style}=\frac{1}{N}\sum_{i,j}\min_{i^{\prime},j^{\prime}}dist(F_{r}(i,j),F_{s}(i^{\prime},j^{\prime})),$ (3)
@@
where $N$ is the number of pixels in $F_{r}$, $dist(a,b)$ is the cosine distance between two feature vectors $a$ and $b$:
$dist(a,b)=1-\frac{a\cdot b}{\|a\|\|b\|}.$ (4)
Method
GT2-GS Framework Overview
In this section, we provide a detailed overview of the proposed GT2-GS framework. Geometry-aware Texture Transfer Loss (GT2 Loss) is proposed to enable controllable and high-quality texture transfer. It is computed using texture features bound with geometric parameters and the prior information from the previous viewpoints. Furthermore, we propose the Adaptive Fine-grained Control Module (AFCM) to regulate the texture learning intensity based on the scene information density in the pixel space. This prevents excessive information from being corrupted by low-granularity features. Finally, for the input Gaussian, an additional color parameter $c^{g}$ is embedded and initialized using the current color parameter $c$. The Geometry Preservation Branch (GPB) is introduced based on $c^{g}$, with the aim of decoupling the optimization processes of the geometry and appearance of the scene. The pipeline is shown in Fig. 2.
@@
Figure 3: Differences in Perspective. The same textured region (highlighted in red) on the 3D object exhibits different texture orientations under varying viewpoints.
Geometry-aware Texture Transfer Loss
Existing methods primarily focus on artistic style transfer and are mostly based on NNFM loss. As shown in Eq. 3, the objective considers only the relationship between the rendered feature map and the style feature map. The optimization objectives across different viewpoints are constructed independently, without accounting for the underlying geometric relationships between views. However, texture is inherently tied to the geometry of the scene. To achieve accurate texture transfer, we incorporate geometric information into the loss function computation during the optimization process.
First, we associate geometric parameters with each texture feature. Considering the perspective geometry within the scene, edge geometric structures, and transformations across viewpoints, we apply scaling and rotation operations to the texture images. Texture features are then extracted from the transformed images and aggregated into a feature set. Specifically, we obtain the original scene depth map using Eq. 2. The depth values are then sorted and discretized into $K$ groups based on predefined depth intervals. The scaling factor for each group is computed as the ratio between the lowest group depth value $Z_{1}$ and the group’s depth value $Z_{k}$. Next, each scaled image is rotated by an angle $\theta$ to capture multi-directional information present in rendered views. Each feature in the texture feature set is denoted as $\{f_{k,\theta}\}$. The scaling parameter $k$ and rotation angle $\theta$ are retained for subsequent computations.
Based on the constructed set of texture features, we propose the GT2 Loss to enable geometry-consistent texture transfer. We construct a per-pixel matched target feature map $F_{t}$ for the rendered feature map $F_{r}$. Texture transfer is achieved by minimizing the cosine similarity between the corresponding feature vectors of $F_{r}$ and $F_{t}$. To ensure that the feature maps $F_{t}$ constructed from multiple viewpoints maintain correct geometric relationships, we incorporate the construction result of the previous viewpoint’s feature map $F_{t}^{v-1}$ when building the target feature map $F_{t}^{v}$ for the current viewpoint. The computation process of the homography matrix $M_{p}^{v,v-1}$ between two viewpoints is formulated as
$M_{p}^{v,v-1}=K_{v-1}[R_{v-1}|T_{v-1}]{[R_{v}|T_{v}]^{-1}}{K_{v}^{-1}},$ (5)
@@
where $K_{v}$ and $R_{v}|T_{v}]$ represent the intrinsic matrix and world-to-camera extrinsic parameters of the $v$-th viewpoint, respectively. Through the homography matrix, the current viewpoint can sample the prior feature $f_{k^{\prime},\theta^{\prime}}$ in the screen coordinate system of the prior viewpoint. Considering occlusion relationships, we filter the projected points using the depth map $I_{d}^{v-1}$ of the prior viewpoint.